there are n teams in a league and the number of games played is given by the polynomial n^2-n=N. A woman's softball league has 7 teams. what is the total number of games to be played?
Answer by ad_alta(240)

If 'N' is the number of games and N(n)=n^2-n, then N(7)=7^2-7=42.